The opera theater manager assumes that 12% of the opera tickets for tonight's show have been sold. If the manager is correct, what is the probability that the proportion of tickets sold in a sample of 473 tickets would be greater than 13%? Round your answer to four decimal places.

Q1

Suppose that a study of elementary school students reports that the mean age at which children begin reading is 5.85.8 years with a standard deviation of 0.80.8 years.

Step 1 of 2:

If a sampling distribution is created using samples of the ages at which 6262 children begin reading, what would be the mean of the sampling distribution of sample means? Round to two decimal places, if necessary.

Step 2 of 2:

If a sampling distribution is created using samples of the ages at which 6262 children begin reading, what would be the standard deviation of the sampling distribution of sample means? Round to two decimal places, if necessary.
Answer #1

Solution :

Given that ,

p = 0.12

1 - p = 0.88

n = 473

= p = 0.12  

= (p*(1-p))/n =  (0.12*0.88)/ 473= 0.01494

a.

P( > 0.13) = 1 - P( < 0.13)

= 1 - P(( - ) / < (0.13-0.12) /0.01494 )

= 1 - P(z < 0.67)

= 1 - 0.7486

= 0.2514

Probability = 0.2514

Q1)

Solution :

Given that,

mean = = 5.8

standard deviation = = 0.80

n = 62

Step 1 of 2:

= = 5.8

Step 2 of 2:

= / n = 0.80/ 62 = 0.10